find four numbers in G.P whose sum is 85 and product is 4096

Grade:10

1 Answers

Ashutosh nayak
7 years ago
See let 4numbers be a,ar^2,ar^3,ar^3.so their sum is a(1+r+r^2+r^3)=85(given) eq 1 .and multiplying them (a^4)*(r^6)=4096 so a*r^1.5=8 eq2 .ok then divide eq 1by eq2it will be (1+r+r^2+r^3)/(r^1.5)=85/8 then by hit and trial get r=4 and putting this value in eq1 get a=1.solved